OpenStudy (amtran_bus):

If 2 and -2 are two zeros of f(x)=x^3-3x^2-4x+12, then the third zero is? A:3 B:4 C:6 D:8 E:12

OpenStudy (amtran_bus):

Is there any way I could use the quadratic formula?

OpenStudy (amtran_bus):

@campbell_st @Jhannybean

OpenStudy (campbell_st):

well I'd factorise it can be done by grouping in pairs \[f(x) = x^3 - 3x^2 - 4x + 12 \] \[f(x) = x^2(x - 3) - 4(x - 3)\] which simplifies to \[f(x) = (x^2 - 4)(x - 3)\] hope this helps

OpenStudy (amtran_bus):

Can you use the quadratic formula at all? Don't you use it for zeros? @campbell_st

OpenStudy (campbell_st):

you could if you had a quadratic factor... since this is a cubic... the quadratic formula only works after finding a linear factor.

OpenStudy (anonymous):

c

OpenStudy (amtran_bus):

Got you.

OpenStudy (amtran_bus):

@maxromeo , it is A

OpenStudy (amtran_bus):

How would you solve it step by step, after what you did campbell?

OpenStudy (jhannybean):

f(x)=(x-3)(x^2-4)=0 solve each part for x (x-3)=0 and (x^2-4)=0 (x-3)=0 x=3 (x^2-4)=0 x^2=4 x= +-2

OpenStudy (amtran_bus):

Thanks